# Count of Monte Crypto: Accounting-based Defenses for Cross-Chain Bridges

### v1

@@ 4.2.2 Other Violating Transactions
PolyNetwork/2021 had one bridge transaction where the withdrawal amount matched the deposit amount, but the withdrawal moved the funds to the wrong destination chain (CrossChecked flagged the mismatch in destination chain specified in the deposit and withdrawal transactions). Anyswap had two groups of bridge transactions in this category. The first consists of 800 withdrawal transactions that pointed to just a few deposits. Manually sampling a few, the deposit transaction hash was not set correctly in the withdrawal transactions (we found their matching deposits), suggesting a program error. The second Anyswap group had four double-spend transactions (referencing the same deposit) that are also likely program errors: the deposit and withdrawal amounts are the same, and the withdrawals occurred within a few hours of each other. Wormhole had two bridge transactions that violated the invariant: 0.5 wrapped Solana on Polygon $\rightarrow$ 650 USDC on Solana, and 4.3k MATIC on Polygon $\rightarrow$ 2.3k on Avalanche. The small amounts and close proximity of the dates of the transactions suggest they are also likely errors.
Finally, three bridges had transactions that had no apparent effect, suggesting invocation errors or undocumented behaviors. PolyNetwork/2023 had one bridge transaction, and Chainswap had four, where the deposits were non-zero but the withdrawal amount was zero. And Anyswap had two withdrawals referencing deposits that did not specify a recipient, yet the deposit amounts covered the withdrawals. While technically not balance invariant violations, CrossChecked flagged them because of their unusual circumstances.
@@
Many of these cases involved highly suspicious unbacked bridge transactions involving very large withdrawals without corresponding deposits. For example, Anyswap had seven transactions totaling more than \$1.5m pointing to non-existent or already-redeemed deposits (e.g., Anyswap #4 and #5 in Table 3 in the appendix). And PolyNetwork/2023 had 27 withdrawals referencing a non-existant deposit address totaling more than \$20m (PolyNetwork/2023 #1 and #2 in Table 3).
@@
The HECO bridge had 73 unusual bridge transactions. One was a very suspicious unbacked bridge transaction minting \$5m of USDT on the HECO chain (HECO #1 in Table 3). The remaining 72, totaling over \$36m, involved withdrawals to an address labeled “HECO recovery”. The label suggests benign intent such as rescuing funds trapped in the bridge, but the activity is also consistent with a rug pull.
@@
Other cases suggest seemingly careless operational practices. In particular, Harmony had 43 bridge transactions that violated the invariant in a variety of different ways. Eight double-spending bridge transactions (e.g., Harmony #1 in Table 3 in the Appendix) used the same deposit to release tokens twice (though only resulting in a profit of a few hundred USD). Thirty-two had indecipherable data: it was not possible to decode the deposit function name, function input, and some events (thus preventing verification of the deposit). One bridge transaction minted piggybankone tokens on Harmony chain referencing a non-existing deposit. And two had withdrawal amounts that were smaller than the deposits, perhaps caused by undocumented fees or errors. Considering the specific mechanism used by the Harmony bridge, where a private key was signing transactions that it should not have, and the fact that some double-spending transactions had amounts different than what was deposited, these incidents suggest that Harmony had issues operating securely and correctly.
5 Live Auditing
Section 4 showed the effectiveness of applying the balance invariant on restropective data to accurately identify past bridge attacks. Incorporating the logic used in the retrospective analysis, we implemented a real-time auditing system that uses the balance invariant to monitor ongoing transactions on a bridge.
@@ 5 Live Auditing
Since most of the bridges in our retrospective analysis shut down at the time of writing (Figure 5), we focus on the Wormhole bridge. Wormhole is still operational, supports a wide variety of blockchains, and is among the most popular bridges in terms of transaction activity [36]. This live system monitors thousands of withdrawal transactions per day across ten blockchains, which together account for 60% of all withdrawal transactions on Wormhole. Extending the system to support other bridges is straightforward.
5.1 System Overview
@@
Figure 6 depicts the architecture of our live auditing system. The system consists of three main components. The Blockchain Monitor tracks the blockchain network and retrieves the latest deposit and withdrawal transactions. The Auditor checks extracted transactions for invariant violations. The Database component stores all the transactions and auditing results.
Figure 6: Live auditing system which consists of three components.
Blockchain Monitor.
@@
The Blockchain Monitor obtains deposit and withdrawal transactions by periodically retrieving the latest finalized blocks. The Monitor uses the same commercial RPC services for collecting Wormhole deposits and withdrawals as we used for the retrospective analysis (Section 4.1). [^12] A fully operational deployment would manage nodes participating in each of the blockchains to obtain transactions in block data directly.
@@
The Monitor only retrieves finalized blocks since unfinalized blocks may be reverted or reordered. Since different blockchains have different finality times, we use the timestamp of the latest finalized block of Ethereum as the synchronization time since Ethereum has the slowest finality time (around 12 minutes) [12]. After retrieving the blocks, the Monitor extracts deposit and withdrawal transactions for Wormhole and saves them to a local database.
The polling interval is configurable and the Monitor polls chains every 1 minute by default. As a result, the live auditing system is as recent as the polling interval combined with the finality time — a situation similar to bridges due to the polling nature of distributed off-chain communication. We leave it to future work to optimize the delay further.
Auditor.
@@
For every new withdrawal transaction extracted by the Blockchain Monitor, the Auditor attempts to find the corresponding deposit transaction in its local database. In most cases it will find the deposit, but if it cannot find it (e.g., because the withdrawal references a non-existent transaction), the Auditor sends an email alert inviting manual inspection. If it finds a deposit, it examines the deposit and withdrawal transactions, and sends an alert if either the deposit has already been withdrawn (the new withdrawal is double-spending) or the bridge transaction violates the balance invariant.
Database.
The system uses a PostgreSQL database to store all information regarding deposit and withdrawal transactions. This information includes the transaction hash, the blockchain, the amount, and the timestamp. It also stores the auditing results, including whether the deposit has been withdrawn and whether the invariant is violated.
5.2 Deployment
We have deployed our live auditing system for the Wormhole bridge for a month. The system audits the deposit and withdrawal transactions between 10 blockchains, which collectively account for around 60% of all withdrawal transactions on Wormhole. On average the observed throughput is approximately one new transaction every two minutes.
The system currently refreshes every one minute, and sends email alerts if it observes a transaction violating the balance invariant or another auditing property (e.g., the destination chain in the deposit does not match the chain in the withdrawal). In the time that we have had the system operational, it has alerted on multiple transactions that violated the invariant. So far these transactions have proven to be errors in our implementation (e.g., failed to extract the correct amount) or the RPC services we used (e.g., not indexing a transaction).
@@ 5.2 Deployment
In addition, while we have not observed any attacks during our monitoring period, we simulated three attack scenarios to confirm that the system alerts when expected. These three transactions represent three types of attacks: (1) a double-spend attack, (2) an unbacked withdrawal attack, and (3) an attack where the deposit and withdrawal amounts do not match. The system successfully alerted on all three attacks.
6 Protecting Bridges
@@
Sections 4 and 5 showed how applying the balance invariant can identify attacks retrospectively and monitor ongoing transactions as an external third party without any changes to existing bridge infrastructure. However, the live monitoring system only alerts on violations of the invariant and does not prevent violating bridge transactions from being processed.
@@
In this section, we describe an approach for extending an existing bridge implementation to prevent malicious transactions from executing. Our goal is to present a proof-of-concept implementation to demonstrate the feasibility of applying the balance invariant in the workflow of bridge transactions as an active defense against malicious transactions. We acknowledge that a number of practical issues remain for a complete operational system, presenting an opportunity for interesting future work.
@@
We start by proposing a new model for the workflow of bridge transactions. Then, using Wormhole as an example, we show how to modify an existing bridge implementation to prevent malicious transactions from being processed with fewer than 30 lines of code changes. Finally, we evaluate the correctness and overhead of this initial implementation.
6.1 The Announce-then-Execute Model
@@
To guide the design of a new model, we first highlight where the attack surfaces exist. As shown in Figure 1, a vulnerability can exist in as early as the second step (verify lock amount) to as late as the next-to-last step (verify relayed transaction). Thus, a natural place to introduce a protection mechanism is just before the final step, which executes the relayed transaction.
Using this insight we propose an announce-then-execute model. Figure 7 depicts the entities and steps involved in the new model, focusing on the destination chain (replacing the right-most box of Figure 1). The new model employs two steps to process a withdrawal transaction: (1) verify and announce the withdrawal transaction on the destination blockchain (steps 9 and 10) and (2) execute the withdrawal transaction (steps 11 and 12). Of particular note, this new model introduces an additional entity, the Approver, which is responsible for approving the relayed transaction before it is executed. The Approver can be a third-party service or a bridge operator, and is responsible for verifying that the announced withdrawal transaction satisfies the invariant by validating the same conditions as the Auditor in the live monitoring system (Section 5). If the relayed transaction violates the invariant, the Approver can reject the transaction to prevent it from being executed.
The announce-then-execute model has a few advantages. First, it explicitly separates the verification and execution of the relayed transaction, thereby providing an opportunity to reject transactions that violate the balance invariant. In comparison, many bridges (e.g., Wormhole) couple the last two steps tightly — the same function call performs both the verification and execution of the relayed transaction. As a result, if the verification step has a bug, it is not possible to stop violating transactions from being executed. Second, the announce-then-execute model allows any external party, that has visibility into both blockchains, to determine if the relayed transaction violates the invariant. A long-standing challenge with bridges is the oracle problem: the destination blockchain does not have visibility into the source blockchain. The announce-then-execute model side-steps the oracle problem by allowing an external party to verify the relayed transaction before it is executed. Moreover, the announce-then-execute model is compatible with existing bridges, as it only requires modification to the last step and treats the relaying infrastructure as well as the verification code as a black box. Lastly, the announce-then-execute model does not introduce any new attack surfaces — even if the approver naively approves all transactions, an attacker still has to exploit other vulnerabilities in the bridge to profit.
Figure 7: Announce-then-execute model for bridges.
6.2 Implementation
Continuing our focus on Wormhole from Section 5, we modified one of Wormhole’s open-source contracts to implement the announce-then-execute model. Wormhole’s current withdrawal function performs two tasks: it verifies the signatures that authorize the withdrawal transaction, and then executes it. We split the withdrawal function into two separate functions: announceWithdraw and approveWithdraw. The announceWithdraw function handles signature verification and announces the withdrawal transaction on the destination blockchain, while the approveWithdraw function executes the withdrawal transaction. This modification required fewer than 30 lines of code changes.
6.3 Evaluation
@@
As a final step we performed a high-level evaluation of the correctness and overhead (in terms of gas usage) of the announce-then-execute implementation. We deployed our bridge implementation on the Binance and Fantom testnets due to the availability of free faucets for acquiring testnet tokens for them. We then executed a series of benign and malicious transactions using Binance as the source chain and Fantom as the destination chain.
@@
Overhead. The additional steps add overhead to implementing bridges. For this experiment, we measured the gas usage for executing Wormhole’s original implementation (with its security checks) and the gas usage of the announce-then-execute implementation (also with the original security checks plus our added code). On average, our proof-of-concept implementation consumed 110k more gas than the original Wormhole implementation (370k vs. 260k gas) when executing the benign transactions, which translates to a roughly \$1 increase for a typical withdrawal on Ethereum using Wormhole.
We note that our implementation is not optimized for gas usage, and the results indicate that a real deployment of the announce-then-execute model will want to optimize the gas overhead of the additional work (e.g., narrowing to transactions above a threshold amount of funds). As discussed further in Section 8, such optimizations and other practical issues of a real deployment are interesting future work.
7 Related Work
@@
Blockchains are, by their nature, public and thus support direct empirical analysis of past transactions. This has engendered a rich literature quantifying and characterizing a range of quasi-adversarial activities on individual blockchains including arbitrage [19, 24], sandwich attacks [24, 39], frontrunning [5], transaction replays [24], and a range of smart contract vulnerabilities and attacks [23, 10, 26, 37, 33, 9]. While our work focuses on the cross-chain context, a number of the vulnerability classes identified in such work are directly implicated in the attacks we analyze.
@@
In the cross-chain context, there are several different streams of related work. First are the efforts to improve the security and performance of cross-chain bridge designs — particularly in managing cross-chain consensus concerning state changes — using zero-knowledge [34] or committees of validators[13, 15]. Some many-chain blockchain ecosystems (e.g., Avalanche) are extending their chains with built-in bridges (across their different chains). We consider our work orthogonal to these efforts as we focus on unbalanced bridge transactions, independent of the particular security violation that allowed such an outcome to take place. Indeed, some of the compromised bridges did use e.g., set of validators.
@@
Another line of work has reviewed real-world attacks on cross chain bridges[14, 36, 21, 38] — ranging from just a few such events to an analysis of over 30 attacks. Our work has directly benefited from the insight and documentation these authors provide, but our respective focus is distinct. While these efforts have concentrated on identifying the vulnerabilities and mechanisms of attack, our work is agonistic to these details and focuses on the financial side-effects those actions.
@@
Yet a third research direction seeks to automatically discover new vulnerabilities in cross-chain bridges using either machine learning (e.g., ChainSniper [29], which trains models to identify vulnerable smart contracts and Lin et al. [17], which train models to detect fake deposit events) or static analysis (e.g., XGuard [30], which statically analyzes bridge contracts for inconsistent behaviors, and SmartAxe [16], which analyzes the control-flow graph of smart contracts and identifies access control and semantic vulnerabilities).
@@
The papers philosophically closest to ours are those that consider the security properties of cross-chain bridges at a higher-level of abstraction. For example, Belichior et al. [2] build and evaluate a synthetic bridge designed to allow high-level monitoring of bridge behavior — including variations in financial state. In a more empirical context, Huang et al. [11] characterize the transactions of the Stargate bridge, and highlight the correlation between large or unusual trades and individual attacks. Finally, perhaps the closest work to our own is Zhang et al.’s XScope [35] which models real-world bridge bugs using a set of pre-defined rules and applies these empirically to identify several attacks retrospectively. Like us, they focused on invariant patterns than the low-level details of smart contract bugs. That said, the rules considered by XScope are still considerably lower-level and more granular than our balance invariant as they are seeking to identify the cause of the detected attacks that have taken place. While two efforts are complementary, we believe our work is simpler to understand and implement, more clearly robust (we have tested against a far wider range of bridges and blockchains) And, as a result, is likely more attractive for inline deployment.
@@
Finally, a range of blockchain companies advertise tools that actively monitor cross-chain bridge transactions for anomalies. These include Hexagate, Hypernative, PeckShield, Slowmist, Certik and ChainAegis (among others), all of which claim various levels of automated alerting for suspicious transactions. However, without clear information on how these systems operate, or empirical results on their efficacy, it is hard to place our work in context beyond our shared goals.
8 Discussion and Future Work

### v3

@@ 5.2.2. Other Violating Transactions
PolyNetwork/2021 had one bridge transaction where the withdrawal amount matched the deposit amount, but the withdrawal moved the funds to the wrong destination chain (CrossChecked flagged the mismatch in destination chain specified in the deposit and withdrawal transactions). Anyswap had 238 withdrawal transactions that pointed to just a few deposits. Manually sampling a few, the deposit transaction hash was not set correctly in the withdrawal transactions (we found their matching deposits), suggesting a program error. In addition, Anyswap had two withdrawals referencing deposits that did not specify a recipient, yet the deposit amounts covered the withdrawals. While technically not balance invariant violations, CrossChecked flagged them because of their unusual circumstances.
Wormhole had two bridge transactions that violated the invariant: 0.5 wrapped Solana on Polygon $\rightarrow$ 650 USDC on Solana, and 4.3k MATIC on Polygon $\rightarrow$ 2.3k on Avalanche. The small amounts and close proximity of the dates of the transactions suggest they are also likely errors.
Finally, three bridges had transactions that had no apparent effect, suggesting invocation errors or undocumented behaviors. PolyNetwork/2023 had one bridge transaction, and Chainswap had four, where the deposits were non-zero but the withdrawal amount was zero.
@@
Many of these cases involved highly suspicious unbacked bridge transactions involving very large withdrawals without corresponding deposits. For example, Anyswap had 29 transactions totaling more than \$1.5m pointing to non-existent or already-redeemed deposits (e.g., Anyswap #4 and #5 in Table 3 in the appendix). PolyNetwork/2023 had 27 withdrawals referencing a non-existent deposit address totaling more than \$20m (PolyNetwork/2023 #1 in Table 3).
@@
The HECO bridge had 73 unusual transactions. One was a very suspicious unbacked bridge transaction minting \$5m of USDT on the HECO chain (HECO #1 in Table 3). The remaining 72, totaling over \$36m, involved withdrawals to an address labeled “HECO recovery”. The label suggests benign intent such as rescuing funds trapped in the bridge, but the activity is also consistent with a rug pull.
@@
Other cases suggest seemingly careless operational practices. In particular, Harmony had 43 bridge transactions that violated the invariant in different ways. Eight double-spending bridge transactions (e.g., Harmony #1 in Table 3) used the same deposit to release tokens twice (though only resulting in a profit of a few hundred USD). Thirty-two had indecipherable data: it was not possible to decode the deposit function name, function input, and some events (thus preventing verification of the deposit). One bridge transaction minted tokens on Harmony chain referencing a non-existent deposit. And two had withdrawal amounts that were smaller than the deposits, perhaps caused by undocumented fees or errors. Considering the specific mechanism used by the Harmony bridge, where a privileged submitter key submitted transactions that it should not have, it suggests that Harmony had issues operating securely and correctly.
6. Live Auditing
@@
Section 5 shows that the balance invariant can accurately identify past bridge attacks with a small amount of manual effort. In this section, we show that this invariant can be used to monitor ongoing transactions in a production bridge in a fully automated fashion. Since most of the bridges in our retrospective analysis are shut down at the time of writing (Figure 4), we focus on the Wormhole bridge. Wormhole is still operational, supports a wide variety of blockchains, and is among the most popular bridges in terms of transaction activity (Zhang et al., 2023). This live system monitors around 1k withdrawal transactions per day across 10 blockchains (Wormhole operates on 10 of the 21 blockchains we support), which roughly account for 40% of all withdrawal transactions. Extending the system to support other bridges is straightforward.
6.1. System Overview
@@
Our system has two main components. The Blockchain Monitor tracks the blockchain network and retrieves the latest deposit and withdrawal transactions. The Auditor checks extracted transactions for invariant violations.
@@
Blockchain Monitor. The Blockchain Monitor obtains deposit and withdrawal transactions by periodically retrieving the latest finalized blocks. The Monitor uses the same commercial RPC services as we used for the retrospective analysis (Section 5.1).
@@
The Monitor only retrieves finalized blocks since unfinalized blocks may be reverted or reordered. Since different blockchains have different finality times, we use the timestamp of the latest finalized block of Ethereum as the synchronization time since Ethereum has the slowest finality time (around 12 minutes) (Konstantopoulos and Buterin, 2021). After retrieving the blocks, the Monitor extracts deposit and withdrawal transactions for Wormhole and saves them to a local database. The Monitor polls every minute (configurable) for new finalized blocks.
@@
Auditor. For every new withdrawal transaction extracted by the Blockchain Monitor, the Auditor attempts to find the corresponding deposit transaction in its local database. In most cases it will find the deposit, but if it cannot find it (e.g., because the withdrawal references a non-existent transaction), the Auditor sends an email alert inviting manual inspection. If it finds a deposit, it examines the deposit and withdrawal transactions. To compute the amount of tokens transferred, the Auditor replays the transactions, which is completely automated and involves no manual effort. The auditor sends an alert if either the deposit has already been withdrawn (the new withdrawal is double-spending) or the bridge transaction violates the balance invariant.
6.2. Deployment
We have deployed our live auditing system for the Wormhole bridge over 10 months (October 2024 to June 2025). The system audits the deposit and withdrawal transactions between 10 blockchains, which we estimate account for around 40% of all withdrawal transactions on Wormhole. Our system audited over 202k transactions (more than 900 transactions per day) and sends email alerts if it observes a transaction violating the balance invariant or another auditing property (e.g., the destination chain in the deposit does not match the chain in the withdrawal).
In the time that our system has been operational, it has alerted on 103 transactions. Upon manual inspection, we confirmed that all alerts were caused by bugs in our code or the RPC services we used (e.g., failing to index a transaction, which typically resolves after retrying after a few minutes). Consistent with our retrospective analysis, the alert rate is very low at 0.05% and raises less than one alert per day.
@@ 6.2. Deployment
In addition, while we have not observed any attacks during our monitoring period, we simulated three attack scenarios to confirm that the system alerts when expected. These three transactions represent three types of attacks: (1) a double-spend attack, (2) an unbacked withdrawal attack, and (3) an attack where the deposit and withdrawal amounts do not match. The system successfully alerted on all three.
7. Protecting Bridges
Figure 5. Report-then-execute model for bridges. The reporter asynchronously reports valid transactions and announces them through events.
@@ 7. Protecting Bridges
Sections 5 and 6 showed how applying the balance invariant could identify attacks retrospectively and monitor ongoing transactions as an external party without any changes to existing bridge infrastructure. However, the live monitoring system only alerts on violations of the invariant and does not prevent violating bridge transactions from being processed.
@@
In this section, we describe an approach for extending an existing bridge implementation to prevent malicious transactions from executing. Our goal is to present a proof-of-concept implementation to demonstrate the feasibility of applying the balance invariant in the workflow of bridge transactions as a defense against malicious transactions.
@@
We start by proposing a new model for the workflow of bridge transactions. Then, using Wormhole as an example, we show how to modify an existing bridge implementation to prevent malicious transactions from being processed with fewer than 40 lines of code changes. Finally, we evaluate the correctness and overhead of this initial implementation.
7.1. The Report-then-Execute Model
@@
To guide the design of a new model, we first highlight where the attack surfaces exist. As shown in Figure 1, a vulnerability can exist in as early as the second step (verify deposit amount) to as late as the next-to-last step (verify relayed transaction). Thus, a natural place to introduce a protection mechanism is just before executing the relayed transaction.
In our first attempt, we explored a simple mechanism that split the submission and execution of the relayed transaction into two steps. Though simple, this mechanism requires the submitter to interact with the bridge twice and increases latency. That said, this earlier iteration, which we included in the preprint version of this paper on arXiv, directly inspired Bascule (Cubist, 2024), a production project that targets a similar but distinct use case (cross-chain transfers between ETH and Bitcoin) and optimizes the process for real-world deployment.
We implement a solution similar to theirs on top of Wormhole and term this solution the report-then-execute model. Figure 5 depicts the entities and steps involved in the new model, focusing on the destination chain (replacing the right-most box of Figure 1). The new model introduces a check step (step 12) before execution. In addition, it introduces a new role, Reporter, which is responsible for asynchronously reporting valid withdrawal transactions. Later, when the submitter submits a withdrawal transaction, the bridge checks with the reporter before execution. The reporter can be a bridge operator, or a set of trusted or trustless third-party (e.g., Bascule (Cubist, 2024) uses trusted execution environments).
In the new model, for a valid withdrawal transaction to execute, the following steps are performed: (1) the reporter asynchronously reports valid transactions that they observe (e.g., to their own contract) and announces transactions they reported through events (steps 8 and 9); (2) the submitter, upon observing that their transaction has been announced, submits the transaction; and (3) the bridge, after processing the transaction (step 11) and before executing it, consults the reporter (e.g., by calling to the reporter’s contract) to verify that the transaction has been reported and the withdrawal amount satisfies the balance invariant (step 12). If the reporter confirms that the transaction is valid, the bridge executes it; otherwise, it rejects the transaction and reverts.
7.2. Assumptions, and Trade-offs
We assume that the reporter can independently verify the relayed transaction and the correct amount of tokens to be transferred. Additionally, we assume the bridge always calls the reporter to verify the transaction before executing it. Under this assumption, the report-then-execute model prevents any implementation bug in any existing bridge components from step 2 (e.g., the bridge fails to compute the correct deposit amount) to step 8 (e.g., the bridge fails to verify the signature of a relayed transaction). Further, if the reporter uses secure hardware or a decentralized third-parties (e.g., with multisig), the model can provide additional protection against insider threats and key theft.
The report-then-execute model has a few advantages. Importantly, it is compatible with existing bridge designs and treats the relaying infrastructure as well as the verification code as a black box (i.e., we make no changes to these components). In addition, the submitter does not have to modify their transaction, and the withdrawal process remains one transaction. The only change is that the transaction has to be reported before submission, which can be done asynchronously and with minimal overhead in delay or latency. As we show later below, our design integrates seamlessly into existing bridge architectures, requires minimal code changes, and only incurs a mild amount of gas overhead.
Furthermore, our model does not introduce any new attack surfaces. Even if the reporter’s key is compromised and reports arbitrary transactions, an attacker still has to exploit other vulnerabilities in the bridge to profit.
Third, the report-then-execute model allows any party (e.g., the bridge operator or a third-party), that has visibility into both blockchains, to determine if the relayed transaction violates the invariant.
Lastly, our model provides an additional opportunity to reject transactions by providing a buffer between the verification and execution steps. In many bridge designs, if the verification step has a bug, it is not possible to stop violating transactions from being executed. In contrast, our model allows the bridge to reject transactions using the simple balance invariant, which is easy to implement and verify.
7.3. Implementation
Continuing our focus on Wormhole from Section 6, we modified Wormhole’s withdrawal function to implement the report-then-execute model. We start by implementing a reporter contract that reports transactions and includes a function for checking if a transaction has been reported. We then modify Wormhole’s contract. Specifically, Wormhole’s current withdrawal function performs two tasks: it verifies the signatures and then executes it. We add a call to the reporter contract such that a verified transaction is checked against the reporter before execution. In total, we made fewer than 40 lines of code changes. For simplicity, we opted for a reporter with one key in our implementation. However, our implementation can be easily extended (e.g., to use multisig).
7.4. Evaluation
@@
As a final step, we performed a high-level evaluation of the correctness and overhead (in terms of gas usage) of the report-then-execute implementation. We deployed our bridge implementation on the Binance and Fantom testnets due to the availability of free faucets for acquiring testnet tokens for them. We then executed a series of benign and malicious transactions using Binance as the source chain and Fantom as the destination chain.
@@
Overhead. The additional steps add overhead to implementing bridges. For this experiment, we measured the gas usage for executing Wormhole’s original implementation (with its security checks) and the gas usage of the report-then-execute implementation (also with the original security checks plus our code). On average, our implementation consumed 120k more gas than the original implementation (380k vs. 270k gas) when executing the benign transactions, which translates to a roughly \$1 increase on Ethereum. We leave optimizations to reduce the gas overhead as future work.
8. Related Work
@@
Blockchains are, by their nature, public and thus support direct empirical analysis of past transactions. This property has engendered a rich literature quantifying and characterizing a range of quasi-adversarial activities on individual blockchains including arbitrage (McLaughlin et al., 2023; Qin et al., 2022), sandwich attacks (Qin et al., 2022; Zhou et al., 2021), frontrunning (Daian et al., 2020), transaction replays (Qin et al., 2022), and a range of smart contract vulnerabilities and attacks (Perez and Livshits, 2021; Grossman et al., 2018; Rodler et al., 2018; Zhang et al., 2020; Wu et al., 2024; Ferreira Torres et al., 2021). While our work focuses on the cross-chain context, a number of the vulnerability classes identified in such work are directly implicated in the attacks we analyze.
@@
In the cross-chain context, there are several different streams of related work. First are the efforts to improve the security and performance of cross-chain bridge designs — particularly in managing cross-chain consensus concerning state changes — using zero-knowledge (Xie et al., 2022) or committees of validators (Lan et al., 2021; Li et al., 2022). Some many-chain blockchain ecosystems (e.g., Avalanche) are extending their chains with built-in bridges (across different chains). We consider our work orthogonal to these efforts as we focus on unbalanced bridge transactions, independent of the particular security violation that allowed such an outcome to take place.
@@
Another line of work has reviewed real-world attacks on cross chain bridges (Lee et al., 2023; Zhang et al., 2023; Notland et al., 2024; Zhao et al., 2023) — ranging from just a few events to an analysis of over 30 attacks. Our work has directly benefited from the insight and documentation these authors provide, but our respective focus is distinct. While these efforts have concentrated on identifying the vulnerabilities and mechanisms of attack, our work is agnostic to these details and focuses on the financial side-effects those actions.
@@
Yet a third research direction seeks to automatically discover new vulnerabilities in cross-chain bridges. Some examples of this work use machine learning such as ChainSniper (Tran et al., 2024), which trains models to identify vulnerable smart contracts, and Lin et al. (Lin et al., 2024), who train models to detect fake deposit events. Other examples use static analysis such as XGuard (Wang et al., 2024) and BridgeGuard (Zhou et al., 2025), which statically analyzes bridge contracts for inconsistent behaviors, and SmartAxe (Liao et al., 2024), which analyzes the control-flow graph of smart contracts and identifies access control vulnerabilities.
@@
The papers philosophically closest to ours are those that consider the security properties of cross-chain bridges at a higher-level of abstraction. For example, Belichior et al. (Belchior et al., 2023) build and evaluate a synthetic bridge designed to allow high-level monitoring of bridge behavior — including variations in financial state. In a more empirical context, Huang et al. (Huang et al., 2024) characterize the transactions of the Stargate bridge, and highlight the correlation between large or unusual trades and individual attacks. Zhang et al.’s XScope (Zhang et al., 2022) and Wu et al. (Wu et al., 2025) model real-world bridge bugs using a set of pre-defined rules and apply these to identify several attacks retrospectively. Compared to our work, the rules they consider are still considerably lower-level and more granular than our balance invariant as they are seeking to identify the cause of the detected attacks that have taken place. Finally, perhaps the closest work to our own is the parallel work by Hu et al. (Hu et al., 2024), which empirically analyzes different aspects of cross-chain bridge transactions. Our work differs in several ways. First, our work is security-oriented and focuses exclusively on token bridges that have been attacked. In comparison, anomaly detection was only a part of Hu et al.’s objectives, while we systematically examine attack surfaces of token bridges. They identified 47 suspicious transactions in PolyNetwork and Anyswap, which is a strict subset of the 80 suspicious transactions we identified. We found more suspicious transactions because we supported a much broader set of blockchains. In addition, we flagged over a hundred more transactions in other bridges that also deserve scrutiny. Most importantly, we propose a simple, generic defense mechanism that can be implemented in-line in existing bridge designs to prevent similar attacks in the future, which differentiates our work from any existing academic work.
@@
Finally, a range of tools exist for monitoring cross-chain bridge transactions for anomalies. These include Hexagate (Hexagate, 2024), Hypernative (Hypernative, 2024), PeckShield (PeckShield, 2024), Certik (CertiK, 2024) and ChainAegis (ChainAegis, 2024) (among others), all of which claim various levels of automated alerting for suspicious transactions. However, without clear information on how these systems operate, or empirical results on their efficacy, it is hard to relate our work to theirs.
9. Discussion and Future Work